Steven Starr has an idea for a desert he is preparing to sell via food truck for restaurant week- grilled bananas with Belgian chocolate ganache. The bananas go through a 4-step process consisting of pealing, filling, grilling, and the last 2 steps are performed simultaneously- splashing and wrapping. Pealing takes 1 minute, filling takes 2 minutes and grilling takes 5 minutes. Splashing takes 3 minute and wrapping takes 3 minutes.

The bottleneck of the process is _____, with an hourly capacity of _____.
A) Grilling, 5
B) Grilling, 12
C) Splashing and wrapping, 10
D) Throughput, 5.45

If Steven Starr’s associate works at restaurant week for a 4-hour flash food festival, what is the maximum number of bananas he can make?
A) 48
B) 20
C) 21.81
D) 17.14

What is the throughput time of the system? _____.
A) 11 minutes
B) 14 minutes
C) 8.5 minutes
D) 5.5 minutes
Answer #1

Answer 1: B: Grilling, 12

Explanation: Processes in last step are done simultaneously, therefore, the bottleneck operation is Grilling with 5 mins process time, the capacity of grilling= 60/process time= 60/5 mins= 12 units per hour

The bottleneck of the process is Grilling with an hourly capacity of 12 units

Answer 2: A)48

Explanation: Production in 4 hours= Hourly capacity*4 hours= 12*4= 48

Answer 3: A)11 mins

Explanation: Throughput time is sum of all process times= Pealing +filling +grilling +splashing = 1+2+5+3= 11 mins